Here is an application problem from a chapter on adding, subtracting and multiplying radicals. Can you help?
Find the exact volume of a cube with the sides of length √3 meters.
thanks.

Answer by stanbon(75887)

Find the exact volume of a cube with the sides of length √3 meters.
thanks.
--------------------------
Volume = side^3
-------
V = (sqrt(3))^3
V = 3^(3/2)
V = 3*sqrt(3)
=============================
cheers,
stan H.